### v1

# The Superclouds of the Local Milky Way

11institutetext: University of Vienna, Department of Astrophysics, Türkenschanzstrasse 17, 1180 Wien, Austria 22institutetext: Istituto di Astrofisica e Planetologia Spaziali (IAPS), INAF, Via Fosso del Cavaliere 100, 00133 Roma, Italy 33institutetext: Max-Planck Institut für Astrophysik, Karl-Schwarzschild-Str. 1, 85748 Garching, Germany 44institutetext: Ludwig-Maximilians-Universität München (LMU), Geschwister-Scholl-Platz 1, 80539 München, Germany 55institutetext: Deutsches Zentrum für Astrophysik, Postplatz 1, 02826 Görlitz, Germany 66institutetext: Excellence Cluster ORIGINS, Boltzmannstr. 2, 85748 Garching, Germany
The Superclouds of the Local Milky Way
Lilly A. Kormann The Superclouds of the Local Milky WayThe Superclouds of the Local Milky Way    João Alves The Superclouds of the Local Milky WayThe Superclouds of the Local Milky Way    Juan D. Soler The Superclouds of the Local Milky WayThe Superclouds of the Local Milky Way    Michelangelo Pantaleoni González The Superclouds of the Local Milky WayThe Superclouds of the Local Milky Way    Cameren Swiggum The Superclouds of the Local Milky WayThe Superclouds of the Local Milky Way    Torsten A. Enßlin The Superclouds of the Local Milky WayThe Superclouds of the Local Milky WayThe Superclouds of the Local Milky WayThe Superclouds of the Local Milky WayThe Superclouds of the Local Milky WayThe Superclouds of the Local Milky WayThe Superclouds of the Local Milky WayThe Superclouds of the Local Milky Way    Gordian Edenhofer The Superclouds of the Local Milky WayThe Superclouds of the Local Milky Way
(Received …; accepted ..)
Recent 3D dust maps of the local Milky Way are revolutionizing our understanding of the Sun’s Galactic neighborhood, bringing much-needed insight into the large-scale organization of the interstellar medium. Focusing on the largest scales in Gaia-based 3D dust maps, we find a pattern of seven highly elongated, mostly parallel structures in the local $5\,\mathrm{kpc}^{2}$, five of which were previously unknown. These structures show pitch angles of $33.5\pm 4.0\degr$, and masses ranging from $10^{5}$ to $10^{6}$ M. We refer to these structures as superclouds. Nearly all known star-forming regions in the solar neighborhood lie within the superclouds, primarily along their central axes, supporting the idea that they are precursors to giant molecular clouds. All but one of the seven superclouds show an underlying undulation, indicating that this is not a property unique to the Radcliffe wave. We find that while the superclouds have line masses that vary by about a factor of four, their volume densities vary by about 10% only. This suggests that superclouds self-regulate their physical sizes and internal structure to maintain pressure equilibrium with their environment. These findings establish a new framework for understanding how large-scale Galactic structures shape the conditions for star formation in the solar vicinity, and likely in galaxies like the Milky Way.
Key Words.:
ISM: clouds – ISM: structure – dust, extinction
1 Introduction
The interstellar medium (ISM) of the Milky Way is a complex and structured environment characterized by features spanning a wide range of spatial scales. Among these, superclouds represent some of the most massive components. The concept of a supercloud first appeared in the seminal work of Elmegreen & Elmegreen (1983), who highlighted the regular distribution of star complexes and their associated neutral atomic hydrogen (Hi) superclouds along galactic spiral arms in nearby galaxies. They proposed that these large-scale structures, with typical masses between $10^{6}$ and $10^{7}$ M and lengths on the order of kiloparsecs, originate from gravitational or magneto-gravitational instabilities in spiral arms. This provided a link between galactic dynamics and the formation of ISM structures, motivating further investigations into the processes that govern the assembly of massive gas concentrations in galaxies.
Expanding on this groundwork, Elmegreen & Elmegreen (1987) identified Hi superclouds in the inner Galaxy and noted their spatial correlation with giant molecular cloud complexes. Similarly, Dame et al. (1986), using carbon monoxide (CO) observations, revealed large molecular complexes that are embedded within superclouds. Grabelsky et al. (1987) conducted a large-scale CO survey of the Carina arm, demonstrating that molecular clouds are additional Population I tracers for this spiral arm, and established a strong link between their distribution and kinematics with those of neutral atomic hydrogen. Lada et al. (1988) presented coordinated HI and CO observations of M31 at the same spatial and spectral resolution to resolve individual giant molecular clouds (GMCs). They find that the M31 GMCs are similar to the Milky Way GMCs and embedded in a large HI spiral arm filamentary feature (see their Figure 1).
Recently, Park et al. (2023) identified several Hi superclouds in the Carina arm associated with H II regions, indicating ongoing star formation within these larger atomic structures, presumably occurring within the embedded molecular clouds. Their analysis of star formation rate surface density also suggests a strong correlation with molecular gas surface density, but little correlation with atomic hydrogen surface density, highlighting the critical role of the molecular phase in the process of star birth.
Subsequent theoretical work on the origin of large-scale ISM structures in galaxies can be found in Franco et al. (2002). They explored the Parker instability in the context of spiral arms, demonstrating via MHD simulations that it can lead to the formation of large-scale gas condensations, thus potentially superclouds. These structures are distributed with a characteristic wavelength along the arm and are preferentially positioned alternately above and below the galactic midplane. Kim & Ostriker (2001, 2002) further emphasized the crucial roles of self-gravity and magnetic fields as the key to gravitational runaway, and explored the formation of large-scale spiral arm features such as spurs, and their fragmentation into GMCs. There is a general consensus from observations and theory that favors a “top-down” (instability), rather than “bottom-up” (coagulation) process for the formation of at least the most massive clouds ($\gtrsim 10^{5}-10^{6}$ M), where most of the molecular material and star formation are found (Kim et al., 2003; McKee & Ostriker, 2007).
Despite their recognition on galactic scales, the properties of superclouds in the immediate solar neighborhood remain poorly constrained. Several factors contribute to this knowledge gap. The isolation of large-scale complexes from the diffuse and extended background of the ISM is a challenge. Identifying individual clouds in Hi is especially difficult due to its widespread and diffuse nature. Combined with line-of-sight overlap, this complicates the establishment of clear structure boundaries. While CO serves as an effective tracer of dense molecular gas, it fails to capture the more diffuse components of the ISM, limiting our ability to trace the full extent of large-scale gaseous structures. Moreover, kinematic distance ambiguities, particularly toward the inner Galaxy, have limited the effectiveness of spectroscopic surveys (e.g., Dame et al., 1986; Elmegreen & Elmegreen, 1987). The possibility that the Sun is embedded within the remnants of an ancient supercloud, as suggested by Olano (2016), further complicates identification. Finally, the definition of a supercloud, whether it emphasizes atomic (Hi) or molecular gas (as traced by CO), varies between studies, making it difficult to establish a uniform and comprehensive survey of such structures in the local Galactic environment.
The recent discovery of the Radcliffe wave, a 2.7 kiloparsec-long chain of dense gas clouds in the solar neighborhood (Alves et al., 2020), has renewed interest in the study of large-scale structures in the ISM. This remarkable structure exhibits an unexpected wavelike shape with a maximum amplitude of about 160 pc, and had remained elusive despite its size and close proximity to the Sun (within 300 pc at its closest point). Like a supercloud, the Radcliffe wave is composed of several massive star-forming regions (SFRs) embedded in well-known GMCs connected by less dense (likely atomic) gas. Its surprising oscillation about the Galactic plane and radial drift away from the Galactic Center (Konietzka et al., 2024) provides a nearby example of a large-scale Galactic-size gaseous feature, but raises questions about the uniqueness of this particular gas arrangement.
To address these challenges, this paper presents a new study of superclouds in the local $\sim 5\,\mathrm{kpc}^{2}$ of the Milky Way. Our approach takes advantage of recent advances in three-dimensional dust mapping from large-scale photometric surveys, particularly those incorporating ESA Gaia data (e.g., Green et al., 2019; Lallement et al., 2019; Chen et al., 2019; Leike et al., 2020; Edenhofer et al., 2024c). By analyzing the 3D distribution of interstellar dust, an effective tracer of total gas column density (e.g., Lada et al., 1994; Alves et al., 1999; Lombardi et al., 2006), we identify regions of enhanced density and find a well-defined pattern of large-scale cloud structures. This method offers a complementary perspective to traditional kinematic analyses and avoids many of the long-standing issues, such as velocity crowding, line-of-sight confusion, and distance uncertainties.
In this paper, we segment the 3D dust map of Edenhofer et al. (2024c), and construct the first catalog of superclouds within the local kiloparsec of the Galaxy. We characterize their physical properties, including filling factor, length, width, height, orientation, and mass, on the basis of the 3D dust distribution, and find a pattern of seven highly elongated, mostly parallel kiloparsec-long superclouds of the local Milky Way. We examine the connection between these superclouds and the broader ISM in the solar neighborhood, and explore potential associations with known stellar populations and other ISM components. Overall, our results provide new insights into the distribution and properties of the most massive interstellar structures in the local Galaxy, with implications for Galactic structure and the early stages of giant molecular cloud formation.
Figure 1: Face-on view of the 40 HOP clouds (left) and their corresponding orientation (right). Each cloud is assigned a distinct color, the length of the line corresponds to the calculated length of the cloud. The Sun is represented by the yellow dot in the center, the Galactic Center is located to the right. An interactive 3D version of the 40 HOP clouds is available here.
2 Data
We use the posterior mean of the 3D Cartesian grid dust map interpolated by Edenhofer et al. (2024c) (hereafter E24), extending out to 1.25 kpc from the Sun. The map is based on stellar distance and extinction estimates from Zhang et al. (2023). To obtain the reconstruction, the logarithm of the dust extinction is modeled using a Gaussian process, implemented in the NIFTy.re framework (Edenhofer et al., 2024b), which incorporates smoothness priors to counteract the ”fingers-of-god” effect. This is enabled by a new generative method called Iterative Charted Refinement (ICR; Edenhofer et al., 2022). ICR efficiently models GPs on nearly arbitrarily spaced voxels, iteratively refining the map’s resolution until a target resolution is reached. The original reconstruction is performed on logarithmically spaced HEALPix spheres with decreasing angular resolution with distance from the Sun. For the posterior mean of the Cartesian interpolated map, we have a resolution of $2^{3}$ pc3, with the data given as unitless extinction values. To convert the data to hydrogen volume density, we apply
$n_{i}=1653\,[\text{cm}^{-3}]\cdot A_{i},$ (1)
where $A_{i}$ is the extinction value in the $i$-th voxel, correlated with hydrogen volume density $n_{i}$ in units of cm-3 via a factor of 1653 cm-3. The derivation of the conversion formula is detailed in O’Neill et al. (2024). By applying this conversion, we obtain a 3D hydrogen volume density distribution map.
3 Methods
To analyze and define the large-scale structures in the dust map, four different methods are tested. The Scikit-Image package (van der Walt et al., 2014), dendrograms, via the astrodendro package (Robitaille et al., 2019), HOP (Colman et al., 2024), a package developed for the extraction of molecular clouds in numerical simulations, and a deep learning approach modified at the Research Center Data Science of the University of Vienna (DS@univie). Ultimately, the HOP algorithm was selected as the best fit for this project. A discussion of the alternative methods and the reasons for their rejection can be found in Appendix A.
3.1 HOP Structure Finder
The structure finding algorithm by Colman et al. (2024) is based on the HOP clump finder of Eisenstein & Hut (1998). HOP was originally designed to identify groups in N-body simulations, and was later adapted by Colman et al. (2024) to extract and compare molecular clouds in numerical simulations. HOP depends on an externally set density threshold and peak density for the data segmentation, which makes a solely data-driven approach impossible. However, it works remarkably well given the large size of the dust map compared to the other methods, and the four segmentation parameters are well-suited for the problem.
3.1.1 Segmentation Stages
HOP goes through two main stages to identify structures, the initial segmentation and the structure definition. Before considering any input parameters, the algorithm checks for density peaks around local volume density maxima, resulting in a preliminary segmentation.
To define a structure in the second stage, the algorithm takes four parameters into account: density_threshold, mincells, density_peak, and density_saddle. The minimum density required for a voxel to be considered as part of a structure is set by density_threshold, the initial density cut. All voxels below this threshold are discarded. The parameter mincells defines the minimum number of cells for an individual structure, structures unable to meet this requirement are discarded or merged with a neighboring structure. To be considered significant, a peak from the initial segmentation must have a minimum density of density_peak. Lastly, density_saddle defines the maximum density threshold for two structures to be seen as individual structures. If two structures have neighboring voxels, HOP identifies boundary pairs, consisting of a border voxel of structure A and structure B. Each boundary pair has a boundary density value that is calculated as the average of their densities. The largest boundary density value per neighboring structures defines the density_saddle value. If the density_saddle value is above the set threshold, the structures are merged and seen as substructures of the main structure. Once the segmentation is done, the field structure_id is added to the data, assigning each voxel to a structure. Voxels with label -1 belong to no structure.
HOP includes a possible third stage to calculate multiple structure properties depending on the available data. This stage is skipped, as the only available data is the density.
3.1.2 Parameter Selection
To find large-scale structures in the E24 dust map, the following HOP parameters are selected after multiple test runs: density_threshold = 1.3 cm-3, density_saddle = 1.3 cm-3, density_peak = 6 cm-3, and mincells = 60000.
The density_threshold was chosen to recover and distinguish the largest present structures, while avoiding merging across the entire map due to low density connections. Increasing density_threshold does not significantly impact the identified clouds, except for slimming them. At higher values, for both density_threshold and density_peak, structures start to disappear, and only the densest regions remain. The parameter density_saddle controls whether structures are separated or merged based on the density of their connecting points. In this work, it is set equal to density_threshold to support the identification of large-scale structures. The purpose of density_saddle is to prevent the connection of structures via thin, low-density bridges, which is not an apparent problem in the final segmentation. Increasing density_saddle results in a more fragmented segmentation, breaking apart coherent clouds, and excluding otherwise present clouds of interest. Lastly, mincells is set to filter out smaller structures that might meet the other criteria, but are not of interest in this study.
Eisenstein & Hut (1998) demonstrated that the density threshold parameter has the greatest influence on the resulting group properties. Changing the initial density cut redefines the outer boundaries of the structures and determines how many voxels are included in the segmentation. In contrast, the remaining parameters primarily influence the internal structure and merging behavior.
3.2 Structure Properties
3.2.1 Mass Calculation
For the mass calculation, we use
$M=1.37\,m_{p}\sum_{i}n_{i}\,dv_{i}$ (2)
from O’Neill et al. (2024). Here, $m_{p}$ is the proton mass, the factor 1.37 is used to convert from hydrogen mass to total mass (Heiderman et al., 2010), and $dv_{i}$ denotes the volume of each voxel. For the Cartesian map, the volume of each voxel is $2^{3}$ pc3.
3.2.2 Shape Definition
To describe the 3D shape, we apply principal component analysis (PCA; Jolliffe 1986 ) to each individual structure. PCA identifies the direction of largest variance within a dataset, determining the primary orientation of the structure. In the case of 3D data, PCA extracts three mutually orthogonal components, creating a new set of basis vectors that aligns with the intrinsic shape of the structure. The first principal component (PC1) describes the direction of the largest variance and the length of the structure. The second principal component (PC2) describes the direction of the largest variance orthogonal to PC1 and the width of the structure. The third principal component (PC3) describes the direction of the largest variance orthogonal to PC1 and PC2, and the height of the structure.
We determine the structure axis lengths by projecting the data onto the principal components and measuring the extent between the minimum and maximum, resulting in an upper boundary. For the length, this is a good approximation, but due to the irregular shape of the clouds, defining the height and width leads to a large overestimation in the majority of the cases.
3.2.3 Orientation Calculation
The structure orientation is determined by the angle between PC1 and a chosen reference axis. This angle is calculated by applying the arccosine to the absolute value of the cosine similarity between the two vectors, and then converting the result to degrees. For consistency and easier comparison, all structures are defined as orientated in the positive direction relative to the reference axis, with angles measured clockwise from it. The angle between PC1 and the Y-axis (face-on) is defined as $\alpha$, while $\beta$ represents the angle between PC1 and the Z-axis (edge-on). Circular statistics are applied to calculate the mean and standard deviation of all angles $\alpha$ and $\beta$. The angles are weighted by mass to consider each structure’s relative importance.
3.2.4 Undulation Fitting
We fit the projected superclouds in the Z/Y plane with a damped sinusoidal of the form
$f(z)=a\,e^{\epsilon_{a}z}\,\sin\left(\omega\cdot e^{\epsilon_{\omega}z}\,z+\phi\right)+d,$ (3)
where $z$ corresponds to the projected Y-coordinate, obtained by rotating the superclouds anticlockwise by their average angle $\alpha$ to lie in the plane, and $f(z)$ gives the corresponding Z-position. The parameter $a$ is the amplitude, $\omega$ the angular frequency, $\phi$ the phase shift, and $d$ the vertical offset. The function includes an exponential modulation of both amplitude and frequency: $\epsilon_{a}$ describes the exponential growth or decay of the amplitude along the wave, while $\epsilon_{\omega}$ describes a linear change in frequency. Positive values for $\epsilon_{a}$ indicate a growth in amplitude over distance, while negative values indicate damping. For $\epsilon_{\omega}$, a positive value indicates a growth in frequency, resulting in a decrease in wavelength via the relation $\lambda=2\pi/\omega$, where $\lambda$ is the wavelength and $\omega$ the angular frequency.
The fitting approach is simplified to give an initial characterization of the underlying undulation, and should not be interpreted as implying a physical oscillation.
4 Results
4.1 HOP Cloud Catalog
HOP identifies 40 clouds in the E24 dust map, visualized in the left panel of Fig. 1. The cloud masses range from $3.8\times 10^{4}$ to $2.8\times 10^{6}$ M, a selection of the calculated cloud properties including the mass and orientation can be found in Table 2.
Figure 2: Orientation distribution of the 40 HOP clouds, the marker color corresponds to the same cloud as in Fig. 1. Distance from the center reflects the cloud mass, the solid line represents the mean orientation. Top: face-on view from the North Galactic Pole, with an average angle of $\alpha=22.6\pm 25.5\degr$. Bottom: edge-on view from the Galactic Center, with an average angle of $\beta=92.3\pm 23.4\degr$.
[…]
The HOP clouds, color-coded by their mean altitude in Z, can be seen in Fig. 3. Larger and more massive clouds are primarily located near the Galactic plane, while smaller clouds are also found at both higher and lower altitudes. Considering the proposed quasi-parallel cloud alignment, underlined by the color gradient in Fig. 1, the color-coding reveals a pattern, where elevated clouds are followed by a downward dip, implying an underlying undulation.
Figure 3: Face-on view of the HOP clouds color-coded by their mean altitude in Z. The Sun is represented by the yellow dot in the center, the Galactic Center is located to the right.
4.2 Supercloud Catalog
Taking the orientation, proximity, and 3D alignment of the HOP clouds into account, we identify seven superclouds, visualized in the left panel of Fig. 4. The corresponding HOP cloud components are listed in Table 1.
To connect the superclouds, the first criterion taken into account is the face-on orientation of the HOP clouds, defined by $\alpha$ and shown in the right panel of Fig. 1. If adjacent clouds show a similar orientation, they are considered as initial candidates for a supercloud. This is verified considering their proximity, candidate clouds that are spatially connected are confirmed as parts of the superclouds. Clouds that deviate from the group orientation, yet still show a spatial connection, are also included as candidate clouds. It is likely that external forces have played a role in the fragmentation of the superclouds during their lifetime. Clouds showing no connection to the supercloud candidates are verified concerning their general 3D alignment. If displaced clouds follow the spatial trend of a larger, already verified group, they are accepted candidates based on their 3D alignment. If these clouds cannot be assigned to another candidate group following the mentioned steps, they are confirmed as part of the first suggested supercloud.
Figure 4: Face-on view of the seven superclouds defined in the local Milky Way (left) and their corresponding orientation (right). Each cloud is assigned a distinct color, the length of the line corresponds to the calculated length of the cloud. The Sun is represented by the yellow dot in the center, the Galactic Center is located to the right. An interactive 3D version of the seven superclouds is available here.
4.2.1 Supercloud Identification
[…]
For the more disrupted orange and yellow cloud group, we adopt the name Malpolon cloud. Its irregular morphology and puffed-up appearance in Z is likely caused by the influence of nearby high-mass stars, as traced by Zari et al. (2023) and the ALS III Catalog of Galatic OB stars; Pantaleoni González et al., under review, which is a recent expansion and Gaia DR3 update of the catalog in Pantaleoni González et al. (2021).
The green supercloud, which we name the Vela Ridge cloud, is the least spatially coherent among the identified superclouds, with its components fragmented and widely separated. A possible explanation is that Vela Ridge itself is not an independent supercloud, but rather a disrupted part of the Sagittarius Spur Extensions, as orbital tracebacks of young star clusters hint that past clustered supernovae may have occurred in this region (Swiggum et al., in prep). This theory of a past connection is also supported by the presence of high-mass stars in the canyon that separates clouds 4 and 33 from cloud 2. The stellar component extends the Spur further to almost double the initially defined length, linking the separated cloud components (Pantaleoni González et al., under review). Spatially, it is not unlikely that clouds 29 and 39 (referred to as the L-clouds) are being sheared off from the SSE due to external forces, such as feedback from high-mass stars in this region (Zari et al. 2023, Pantaleoni González et al., under review), or supernovae from the Messier 6 cluster family that partially occupies this region (Swiggum et al., 2024). Additionally, clouds 23 and 26 may have originally been part of the Split, which is supported by their physical connection. The displacement is likely a result of the event responsible for the ring they surround. We keep Vela Ridge as a supercloud in this work, as we cannot verify any of the alternative interpretations yet.
Lastly, the brown supercloud, located at the upper edge of the map, is given the name Anguis cloud. While the two supercloud components are not connected, comparing our results with the extended version of the Edenhofer et al. (2024c) dust map and the dust map by Vergely et al. (2022) shows that they are likely part of a supercloud extending beyond the map.
4.2.2 Supercloud Properties
All superclouds we identify are truncated by the artificial boundary of the 3D reconstruction and are therefore incomplete. Their cloud components are either cut off at the border of the map, or show potential extensions when compared to the 2 kpc dust map by Edenhofer et al. (2024c) and the dust map by Vergely et al. (2022). These maps are not used for the segmentation itself, due to their decreasing reliability beyond the boundary of the main map and their lower resolution. Instead, they serve as additional references for assumptions on the possible continuation of the superclouds.
As shown by Alves et al. (2020), the Radcliffe wave extends to the Cygnus X complex. Zucker et al. (2020) calculated accurate distances for SFRs in the Star Formation Handbook (Reipurth, 2008a, b) up to 2.5 kpc from the Sun. Comparing their results with the superclouds, we find that the Natrix cloud is likely to extend to the Rosette Nebula, while the SSE connects to M20, Lagoon, L291, NGC6604, M17, M16, and L379. It is interesting to note that each supercloud, excluding the Anguis cloud, contains at least a few known star-forming regions. This absence could indicate a less active or earlier phase of cloud evolution, but it might also result from observational restraints or incomplete coverage of the region.
The masses of the inferred superclouds range from $1.1$ to $3.6\times 10^{6}$ M, with the Anguis cloud as an outlier at $7.8\times 10^{5}$ M. All superclouds, except the Anguis cloud, show lengths ranging from approximately 1.5 to 2.5 kpc, which is calculated using PCA and taking all cloud components into account. A selection of the calculated cloud properties including mass and orientation can be found in Table 3. Their mean orientation is characterized by a face-on pitch angle $\alpha=33.5\pm 4.0\degr$, and edge-on angle $\beta=87.7\pm 4.2\degr$, making an association with a major spiral arm unlikely (Franco et al., 2002; Efremov, 2009). We show the face-on orientation of the superclouds in the right panel of Fig. 4. The identified superclouds are consistent with the characteristic scales and masses predicted in earlier studies (Elmegreen & Elmegreen, 1987; Park et al., 2023), representing the largest scale of ISM structures in the local Milky Way and likely throughout the Galaxy as a whole.
In total, the E24 dust map has a gas mass of $3.3\times 10^{7}$ M. For Z = ± $\pm\,100$ pc, the volume filling factor of the superclouds is 9.94%, containing 55.95% of the gas mass in the Galactic plane.
4.2.3 Undulation Pattern
Notably, the undulation observed in the Radcliffe wave is not unique. All but one of the seven superclouds show an undulating pattern in the Z/Y plane, indicated by the color pattern of Fig. 3, and are likely oscillating (Konietzka et al., 2024). Here, the upper part of the Radcliffe wave appears flat in Z, as a result of averaging the change in height across the entire length of HOP cloud 0, erasing the underlying pattern.
In Fig. 5 we show the sinusoidal fit for the selected superclouds: the Malpolon cloud, Natrix cloud, Radcliffe wave, and Vela Ridge cloud. The SSE and the Anguis cloud are excluded from the analysis, as they show a vertical displacement, but it is not as prominent as for the previous examples. The Split is also excluded, as it appears unusually flat, with only two noticeable downward extensions at Y = -200 pc (referred to as the ”C” and discussed in Edenhofer et al. 2024a) and Y = 500 pc. All resulting fit parameters are listed in Table 4, the fitting approach is simplified to give an initial characterization of the underlying undulation.
For the Malpolon cloud, we exclude the three highest altitude clouds (cloud 10, 16, and 17) while fitting. It is likely that these clouds were pushed up by nearby high-mass stars (Zari et al. 2023, Pantaleoni González et al., submitted to MNRAS), and therefore distort the inherent structure of the supercloud. The Malpolon cloud, Natrix cloud, and Radcliffe wave are well described by a damped sinusoidal wave. For the Malpolon cloud, the parameters indicate decay in the amplitude and growth in the wavelength. Both damping parameters for the Natrix cloud indicate growth over distance. For the Radcliffe wave, the parameters show decay in amplitude and wavelength.
Vela Ridge is fitted with a simple sinusoidal function. While the undulation is visible, the large wavelength is likely a consequence of the fragmented nature of the supercloud. This, together with the incompleteness, support the idea that Vela Ridge may not be an individual supercloud.
Figure 5: 2D projection of selected superclouds in the Z/Y plane with a sinusoidal fit. For the fit, the clouds are rotated anticlockwise by their $\alpha$ in the X/Y plane. The first three clouds are fit with a damped sinusoidal, the fourth cloud with a simple sinus function.
5 Discussion
A key implication of the main results in this paper is that these superclouds are unlikely to be associated with the spiral arms of the Milky Way, which have significantly smaller pitch angles (on average 10, as shown in Reid et al. 2019). This includes the most massive of the local superclouds, the Radcliffe wave, normally associated with the Local Arm (also known as the Orion Arm or the Orion Spur), e.g., Swiggum et al. (2022). Instead, the local superclouds seem to be part of a distinct, coherent structure that may reflect larger-scale dynamical processes in the Milky Way, perhaps akin to the high-pitch-angle gas lanes observed in other galaxies.
Additionally, given that none of the seven superclouds seems to be associated with a Milky Way spiral arm, this implies that there is no obvious external trigger for the formation of GMCs inside them, like a spiral shock. An initial comparison with the spatial distribution of known star-forming GMCs shows that nearly all lie within these superclouds, with many concentrated along their central regions of higher density, a result aligned with earlier results from Dame et al. (1986); Elmegreen & Elmegreen (1987). This result supports a molecular cloud formation scenario in the local Milky Way in which H2 forms in the denser regions of the neutral atomic cold medium without an obvious external trigger mechanism. This is tentative at this point, and more work is needed to qualify this statement. Since GMCs are the primary sites of star formation, the superclouds presented in this work are likely to play an indirect but critical role in shaping the star formation process in the local Milky Way.
Of the seven superclouds identified in this study, only two were previously known: the Split (Lallement et al., 2019) and the Radcliffe wave (Alves et al., 2020). While the Split lies on the Galactic midplane, the Radcliffe wave surprisingly undulates above and below it. Recent work by Konietzka et al. (2024), using young stellar clusters, confirmed that the undulations of the Radcliffe wave represent vertical oscillations relative to the plane. Our results show that such undulations are not unique to the Radcliffe wave. In fact, all superclouds, except the Split, exhibit similar vertical displacements. The origin of these undulations remains uncertain. One possibility is that they could result from shocks or gravitational perturbations as superclouds interact with spiral arms in the past, but further investigation is needed to test this scenario.
The identified superclouds can be associated with multiple known structures in the solar neighborhood. Spatially, the Anguis and the Malpolon cloud appear to be a small section of the Cepheus spur (Pantaleoni González et al., 2021). This Spur extends from Cygnus OB2 towards the Galactic anti-center, as traced by young stellar populations and in more extended dust maps (e.g., Lallement et al., 2019). The bottom part of the Radcliffe wave (cloud 27, 34, 30, 12, and 18) and of the Vela Ridge cloud (cloud 21 and 2) define the boundaries of the GSH 238+00+09 supershell (Heiles, 1998). This kiloparsec long dust shell likely formed from clustered supernovae associated with the Cr135 cluster family (Swiggum et al., 2024), and occupies the full width between the two superclouds. Recently, Soler et al. (2025) calculated the kinetic energy densities of the local ISM based on the E24 dust map. Kinetic energy overdensities represent areas that differ roughly by a factor of ten or more from the mean value. The largest of these overdensities can be connected to the Radcliffe wave and the Split, likely as a result of past supernovae or large-scale galactic dynamics.
Figures 5 and 6 reveal an intriguing clue about the stability of local superclouds. While their mass-per-length varies by factors of 3–4, their average densities differ by only about 10%. This points to a state of approximate pressure equilibrium between the superclouds and the surrounding medium. In such a scenario, external pressure sets the density scale, while individual superclouds adjust their masses and sizes within this constraint, resulting in structurally diverse clouds with similar internal conditions. In other words, although their column densities (and thus gravitational binding) differ, their internal pressures remain roughly comparable. We propose that these structures self-regulate their physical sizes and internal kinematics to maintain a pressure balance with their environment. This would naturally explain the coexistence of superclouds with varying masses and dimensions within a common large-scale ISM context, consistent with observations.
Figure 6: Properties calculated for the seven superclouds. The color corresponds to the same cloud as in Fig. 4. RadWave refers to the Radcliffe wave, SSE to the Sagittarius Spur Extension. Top: total mass of the superclouds. Second: total length of the superclouds. Third: linear mass of the superclouds, the ratio of total mass to total length. Bottom: average density of the superclouds. The trends represent a lower boundary, as all superclouds are incomplete and likely to continue beyond the map.
The well-defined, quasi-parallel pattern of superclouds shown in Fig. 4 suggests a relatively undisturbed local ISM. While stellar feedback is clearly reshaping the gas on kiloparsec scales, for example, in the case of the Heiles supershell (Swiggum et al., 2024), it does not appear to have significantly disrupted the overall supercloud pattern. Based on the formation scenario for GMCs discussed above, we can make a conservative, lower age estimation by considering the ages of the oldest stars formed within the superclouds. For the Radcliffe wave, this corresponds to approximately 20-30 Myr. Compared to kinematic timescales, this lower age estimate is short. Recent measurements of the local ISM velocity dispersion, approximately $11\,\mathrm{km\,s^{-1}}$ (see Soler et al., 2025, and references therein), suggest that a perturbation would require over $90\,\mathrm{Myr}$ to cross the full length of a supercloud. This timescale is about three times longer than the estimate based on stellar ages and exceeds the typical lifetimes of CO clouds (see, for example Miville-Deschênes et al., 2017; Chevance et al., 2020). These discrepancies imply that superclouds are unlikely to form by a bottom-up process, i.e., from small to large scales, as would be expected in a scenario driven purely by stellar feedback. Instead, large-scale Galactic dynamics and instabilities probably play a central role in the formation of the kiloparsec-scale density structures we identify as superclouds (see, for example Kim & Ostriker, 2002, 2017; Smith et al., 2020; Colman et al., 2022). This does not exclude the influence of energy and momentum injection from interior star formation (Swiggum et al., 2024), but it points to the need for additional physical processes to explain the observed coherence and scale of these structures.
6 Conclusion
The most significant outcome of this study is the discovery of a remarkably organized gas distribution within the local kiloparsec of the Milky Way. This distribution follows a distinct pattern of quasi-parallel gas lanes extending over several kiloparsec (see right panel of Fig. 4). These lanes contain substantial masses of the order of 106 M, and exhibit pitch angles of approximately 30. Given their size, mass, and likely composition (predominantly atomic hydrogen), we refer to these structures as superclouds, following the terminology of Elmegreen & Elmegreen (1983). This identification has only become possible with the advent of large-scale 3D dust maps based on data from the Gaia mission (Gaia Collaboration et al., 2023), allowing us to trace the local gas distribution in three dimensions with unprecedented detail. Unlike earlier studies of superclouds relying on Position-Position-Velocity analysis, these new maps offer a clearer, spatially resolved view of the interstellar medium.
The discovery of a coherent network of superclouds within the local kiloparsec marks a step forward in our understanding of the Milky Way’s ISM. These massive, elongated structures appear to serve as scaffolds for molecular cloud formation and star formation on Galactic scales. Much remains to be understood: what these structures truly are, how they form, and how they give rise to GMCs and, ultimately, stars. Their morphology, stability, and apparent independence from the traditional spiral arm framework raise important questions about the large-scale dynamics of the Galaxy. As new data and analysis tools continue to improve our view of the three-dimensional ISM, these findings offer a foundation for future investigations into the interplay between galactic structure, gas dynamics, and star formation. As we refine our three-dimensional view of the nearby Galaxy, we are beginning to define a solid benchmark for connecting the complex structure of the Milky Way with the broader context of galaxy evolution and star formation in the universe.
Acknowledgements.
The authors thank Tine Colman for helpful comments on the HOP algorithm, and Alvaro Hacar for insightful discussions on the topic of superclouds. Co-funded by the European Union (ERC, ISM-FLOW, 101055318). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The results in this paper were based on observations obtained with Gaia, ESA science missions with instruments and contributions directly funded by ESA Member States, NASA, and Canada. This research has made use of NASA’s Astrophysics Data System, the SIMBAD database, and the VizieR catalog access tool operated at CDS, Strasbourg, France, as well as publicly available software libraries, including Astropy (Astropy Collaboration et al., 2013), NumPy (Harris et al., 2020), scikit-learn (Pedregosa et al., 2011), matplotlib (Hunter, 2007), and plotly (Inc., 2015). The computational results have been achieved using the Austrian Scientific Computing (ASC) infrastructure.
Appendix A Rejected Segmentation Methods
Here, the additional three segmentation methods are briefly discussed, along with the reasons for their rejection.
A.1 Scikit-Image
Scikit-Image (van der Walt et al. 2014) is a tool for processing and analyzing 3D data sets, making it possible to extract higher density features in the dust data and labeling connected regions. It offers a fast and straightforward segmentation approach, but has a few limitations that make it unsuitable for this work. First, the method requires the minimum density threshold, the only parameter the segmentation depends on, to be predefined. The data is then segmented by distinguishing between high- and low-density voxels, depending on the given threshold. Although the labeling option makes it possible to identify connected structures, the approach is too simple to allow for any fine-tuning.
A.2 Dendrograms
Astrodendro (Robitaille et al. 2019) is a package for constructing dendrograms; hierarchical structures used to identify and analyze nested features. While dendrograms are effective for analyzing and identifying hierarchical substructures, the size of the dust map makes it difficult to analyze regions of interest, due to the large number of resulting branches. This can be avoided by fine-tuning the parameters, but the more effective way is to extract smaller regions for the analysis. However, turning the map into smaller segments is not the preferred approach.
A.3 Convolutional Neural Network
The convolutional neural network (CNN) approach is based on Kim et al. (2020), who developed the algorithm for unsupervised image segmentation in 2D. It was later modified at the DS@univie (private communication with Amanda de Paula Cristino Hirschl) for 3D applications. Originally adapted to find gaps in the dust map by Leike et al. (2020), it can also be applied to detect high-density regions. In this approach, the CNN processes the dust data by extracting structural features based on density. This is done through convolutional layers, which enable the network to learn spatial patterns on different scales. The output is a binary dataset that indicates whether a data point belongs to a structure or is rejected as part of the background. The CNN enables a data-driven segmentation that does not depend on an external density threshold definition, but is not suitable to further segment the defined large-scale structures. This limitation makes it difficult to properly segment the dust map.
Appendix B Cloud Numbering
Each cloud and supercloud is assigned a number for identification purposes. The HOP cloud numbers are assigned in Fig. 7, the supercloud numbers in Fig. 8. Table 1 indicates which HOP clouds are part of each supercloud.
Figure 7: Same as Fig. 1 (left), with additional cloud ID numbers.
Figure 8: Same as Fig. 4 (left), with additional supercloud ID numbers. ID 0 is the Racliffe wave, 1 th
Table 1: HOP cloud components for each supercloud. The ID refers to the cloud ID as shown in Fig. 7.
ID Supercloud HOP Cloud
0 Radcliffe Wave 0, 12, 18, 27, 30, 34, 37
1 Split 3, 7, 8
2 Natrix Cloud 6, 14, 22, 28, 31, 32, 35
3 Malpolon Cloud 9, 10, 13, 15, 16, 17, 20
25, 19, 24, 36, 38
4 Vela Ridge Cloud 2, 21, 23, 26, 29, 39
5 Sagittarius Spur Extension 1, 4, 33
6 Anguis Cloud 5, 11
Appendix C Cloud Properties
Multiple properties are calculated for the HOP clouds as described in Sect. 3 and can be found in Table 2. The supercloud properties are shown in Table 3. The resulting fit parameters for the Malpolon cloud, Natrix cloud, Radcliffe wave and the Vela Ridge cloud are shown in Table 4.
Table 2: Selection of parameters calculated for the 40 HOP clouds. Each cloud is assigned a cloud ID and has listed which supercloud it is part of. The column cell count indicates the voxels per structure, length (L) gives the cloud length calculated with PCA. The columns X, Y, and Z are the unweighted mean coordinates per cloud. Alpha $\alpha$ and beta $\beta$ represent the face-on and edge-on angles of the clouds.
[…]
   0 Radcliffe Wave 3142543 2.77e+06 1429.77 114.76 719.92 -15.75 31.46 91.94
   1 Sag. Spur Extension 2018035 1.81e+06 1100.18 1035.86 -289.03 -2.24 21.24 91.37
   2 Vela Ridge Cloud 1168440 1.06e+06 654.03 -54.58 -893.01 18.71 31.39 101.79
   3 Split 990721 9.39e+05 742.11 426.96 383.43 19.78 10.11 89.68
   4 Sag. Spur Extension 693664 5.24e+05 466.71 683.74 -580.29 39.30 298.92 144.88
   5 Anguis Cloud 675154 5.05e+05 495.69 -928.40 655.21 19.14 30.51 50.88
   6 Natrix Cloud 569140 5.14e+05 543.46 -311.49 722.62 60.55 10.18 78.85
   7 Split 535072 5.43e+05 554.04 189.72 18.26 11.00 30.61 75.75
   8 Split 474359 3.67e+05 466.99 727.94 683.38 0.83 30.37 89.98
   9 Malpolon Cloud 415170 3.06e+05 413.61 -564.91 726.98 35.75 13.46 133.17
   10 Malpolon Cloud 376439 3.17e+05 461.37 -306.69 999.70 243.40 9.10 102.14
   11 Anguis Cloud 343470 2.71e+05 435.42 -601.23 1014.20 33.62 41.08 37.51
   12 Radcliffe Wave 338392 2.93e+05 504.32 -698.04 -529.46 -179.72 43.88 112.12
   13 Malpolon Cloud 318522 2.32e+05 333.37 -723.67 566.25 54.48 25.75 68.03
   14 Natrix Cloud 306527 2.47e+05 516.19 -449.49 227.49 -30.37 12.84 81.06
   15 Malpolon Cloud 302052 2.19e+05 524.56 -722.79 315.55 66.79 67.97 112.46
   16 Malpolon Cloud 277775 1.64e+05 323.23 -784.05 552.04 167.18 44.67 85.16
   17 Malpolon Cloud 247255 1.66e+05 310.93 -652.01 722.71 176.71 328.35 65.37
   18 Radcliffe Wave 247032 2.82e+05 369.44 -397.34 -177.36 -103.16 14.40 77.41
   19 Malpolon Cloud 188997 1.40e+05 411.70 -1019.06 -254.26 -23.77 27.12 88.54
   20 Malpolon Cloud 188243 1.43e+05 454.32 -1032.00 10.69 88.33 13.44 90.00
   21 Vela Ridge Cloud 159910 1.09e+05 315.85 -334.22 -997.13 -70.90 85.58 90.44
   22 Natrix Cloud 159249 1.26e+05 279.32 -166.65 909.66 153.50 48.65 131.09
   23 Vela Ridge Cloud 145303 8.44e+04 283.46 819.62 335.96 137.91 345.86 137.22
   24 Malpolon Cloud 138999 1.06e+05 319.05 -422.34 1066.74 107.69 62.70 160.71
   25 Malpolon Cloud 138713 1.13e+05 295.11 -453.36 903.42 47.68 281.21 8.52
   26 Vela Ridge Cloud 133466 9.65e+04 220.03 818.50 523.70 86.36 310.51 147.82
   27 Radcliffe Wave 115059 8.50e+04 213.97 -862.93 -814.00 -66.09 68.35 166.31
   28 Natrix Cloud 104836 8.50e+04 266.80 -725.34 -272.38 34.74 56.11 113.74
   29 Vela Ridge Cloud 104127 1.02e+05 262.17 749.41 37.72 95.62 306.39 93.83
   30 Radcliffe Wave 95155 6.65e+04 231.40 -638.95 -714.05 -131.42 37.42 114.98
   31 Natrix Cloud 85936 5.87e+04 236.84 -684.91 -145.18 -5.22 81.38 33.35
   32 Natrix Cloud 84333 5.50e+04 236.98 -1021.12 -613.86 14.61 32.79 36.12
   33 Sag. Spur Extension 72527 5.22e+04 271.71 546.04 -964.51 126.08 1.19 80.83
   34 Radcliffe Wave 67317 4.77e+04 275.54 -766.10 -860.39 -154.85 355.28 92.10
   35 Natrix Cloud 63846 3.77e+04 192.22 -915.17 -497.46 -51.00 17.88 96.27
   36 Malpolon Cloud 62484 4.58e+04 220.50 -664.41 612.74 -34.04 27.61 71.41
   37 Radcliffe Wave 61628 4.13e+04 176.90 -324.72 308.98 82.87 45.88 100.38
   38 Malpolon Cloud 61550 3.85e+04 176.38 -1097.81 -124.34 0.56 345.45 109.98
   39 Vela Ridge Cloud 60876 4.64e+04 282.88 565.15 -277.92 113.30 32.39 103.70
Table 3: Selection of parameters calculated for the seven superclouds. Each supercloud is assigned a supercloud ID and name. The column cell count indicates the voxels per structure, length (L) gives the cloud length calculated with PCA. The columns X, Y, and Z are the unweighted mean coordinates per cloud. Alpha $\alpha$ and beta $\beta$ represent the face-on and edge-on angles of the clouds.
[…]
   Radcliffe Wave 4067126 3.58E+06 2468.67 -50.50 452.14 -39.64 31.42 86.67
   Split 2000152 1.85E+06 1399.60 434.87 356.88 12.94 36.87 90.02
   Natrix Cloud 1373867 1.12E+06 1970.08 -452.04 364.88 36.95 25.67 85.19
   Malpolon Cloud 2716199 1.99E+06 1946.38 -660.98 568.46 97.84 30.02 82.99
   Vela Ridge Cloud 1772122 1.50E+06 2215.88 126.15 -619.12 33.26 36.34 85.31
   Sagittarius Spur Extension 2784226 2.39E+06 1487.87 935.37 -379.19 11.45 35.85 95.50
   Anguis Cloud 1018624 7.76E+05 981.80 -818.08 776.26 24.02 42.19 82.99
Table 4: Fit parameters for the undulating superclouds. The column $a$ denotes the amplitude, $\lambda$ the wavelength, $\phi$ the phase shift, and $d$ the vertical offset. The exponential damping parameters $\epsilon_{a}$ for the amplitude and $\epsilon_{\omega}$ for the frequency are only listed for clouds fitted with a damped sinusoidal, as described in Sect. 3.2.4, instead of a standard sinusoidal function.
  Supercloud Amplitude $a$ [pc] Wavelength $\lambda$ [pc] Phase $\phi$ [rad] Offset $d$ [pc] Damp. $\epsilon_{a}$ [pc-1] Damp. $\epsilon_{\omega}$ [pc-1]
   Malpolon Cloud 31.97 1137.40 -3.10 58.15 $-1.18\times 10^{-3}$ $-7.52\times 10^{-5}$
   Natrix Cloud 70.53 1749.47 -0.58 -1.17 $1.59\times 10^{-3}$ $-1.12\times 10^{-3}$
   Radcliffe Wave -87.95 1748.00 -3.09 -13.01 $-1.10\times 10^{-3}$ $2.88\times 10^{-4}$
   Vela Ridge Cloud 56.90 3078.95 0.41 59.14